### v1

# Unlocking Multimodal Integration in EHRs: A Prompt Learning Framework for Language and Time Series Fusion

Abstract
Large language models (LLMs) have shown remarkable performance in vision-language tasks, but their application in the medical field remains underexplored, particularly for integrating structured time series data with unstructured clinical notes. In clinical practice, dynamic time series data such as lab test results capture critical temporal patterns, while clinical notes provide rich semantic context. Merging these modalities is challenging due to the inherent differences between continuous signals and discrete text. To bridge this gap, we introduce ProMedTS, a novel self-supervised multimodal framework that employs prompt-guided learning to unify these heterogeneous data types. Our approach leverages lightweight anomaly detection to generate anomaly captions that serve as prompts, guiding the encoding of raw time series data into informative embeddings. These embeddings are aligned with textual representations in a shared latent space, preserving fine-grained temporal nuances alongside semantic insights. Furthermore, our framework incorporates tailored self-supervised objectives to enhance both intra- and inter-modal alignment. We evaluate ProMedTS on disease diagnosis tasks using real-world datasets, and the results demonstrate that our method consistently outperforms state-of-the-art approaches.
Unlocking Multimodal Integration in EHRs: A Prompt Learning Framework for Language and Time Series Fusion
Shuai Niu1, Jing Ma1, Hongzhan Lin1, Liang Bai2, Zhihua Wang3, Wei Bi5, Yida Xu1, Guo Li4, and Xian Yang6thanks: * This is the corresponding author. Hong Kong Baptist University1, Shanxi University2, Shanghai Institute for Advanced Study of Zhejiang University3, Manchester Metropolitan University4, Tencent AI Lab5, The University of Manchester6 {cssniu,majing, cshzlin}@comp.hkbu.edu.hk, sxbailiang@126.com, zhihua.wang@zju.edu.cn, victoriabi@tencent.com, xuyida@hkbu.edu.hk, l.guo@mmu.ac.uk, xian.yang@manchester.ac.uk
Figure 1: (a) LLMs struggle to process continuous time series data due to modality gaps with discrete textual representations. (b) ProMedTS bridges this gap by leveraging anomaly descriptions and time series prompts, aligning structured EHR data with clinical notes for improved multimodal understanding.
1 Introduction
Recent advancements in natural language processing (NLP) have revolutionized healthcare by enabling deeper insights into electronic health records (EHRs). EHRs combine structured data, such as time series laboratory (lab) test results, with unstructured data, including clinical notes and medical images. While large language models (LLMs) excel at processing unstructured text Nori et al. (2023); Singhal et al. (2023) and vision transformers have driven progress in medical image analysis Wang et al. (2022); Chen et al. (2021), integrating continuous time series data with text remains a challenge. Unlike text, which is composed of discrete tokens, time series data contain continuous signals with temporal dependencies Jin et al. (2023) as illustrated in Figure 1(a).
Current multimodal learning approaches, especially contrastive learning methods Radford et al. (2021); Li et al. (2023), have been effective in aligning vision and language. However, they are less suited to bridge the gap between time series and text. Time series data require fine-grained temporal representations in a high-dimensional space and are often irregularly sampled, exhibit diverse frequencies, and include missing values Harutyunyan et al. (2019a). In addition, the lack of large-scale paired datasets that link raw time series with textual descriptions further hampers LLMs from incorporating structured information into clinical decision-making Niu et al. (2024). Without an effective fusion mechanism, LLMs cannot fully exploit the rich temporal patterns in structured EHR data.
To address these challenges, we propose ProMedTS, a self-supervised and prompt-guided framework designed to unify medical notes and time series lab test for naturally understood by LLMs. As shown in Figure 1(b), instead of feeding raw time series directly into LLMs, our framework uses anomaly descriptions, capturing key patterns in lab test results for helping multimodal fusion between text and time series data. These descriptions are generated using lightweight anomaly detection techniques Vinutha et al. (2018) and convert continuous signals into human-readable summaries. The process involves two steps. First, anomaly descriptions establish a direct connection between time series EHRs and medical notes. Second, time series prompt embeddings are generated and added as prefix tokens to the LLM input. This method integrates structured time series information into the language modeling process without altering the LLM architecture, unifying both modalities within a same encoding space and enhancing clinical decision-making.
We optimize ProMedTS with three self-supervised learning objectives. A contrastive loss maps textual and time series modalities into a shared latent space. An anomaly-time series matching loss links lab test with their corresponding anomaly descriptions to reinforce consistency. Finally, an anomaly caption generation loss improves the fine-grained alignment between numeric time series lab test and time series prompt embeddings. Together, these objectives enable LLMs to process both structured and unstructured EHR data more effectively, addressing the gap between language and time series representations in healthcare applications.
  • We propose ProMedTS, a self-supervised framework that integrates structured time series and unstructured textual EHR data into LLMs without changing their architectures.
  • We introduce anomaly descriptions as a textual bridge to align time series data with clinical notes, supported by three self-supervised objectives.
  • We demonstrate that ProMedTS significantly improves disease diagnosis on MIMIC-III and MIMIC-IV, setting a new benchmark for multimodal EHR learning.
2 Related Work
The increasing diversity of EHR data has led to significant advancements in multimodal learning for healthcare applications. MedCLIP Wang et al. (2022) employs semantic contrastive learning to align medical images with textual reports, while RAIM Qiao et al. (2019) and GLoRIA Huang et al. (2021) integrate numerical and image data with text using attention mechanisms. LDAM Niu et al. (2021a) further extends these approaches by leveraging cross-attention with disease labels to fuse features from lab tests and clinical notes. EHR-KnowGen Niu et al. (2024) transforms structured lab data into text and incorporates external knowledge for improved modality fusion. Despite these advancements, achieving a unified latent embedding that effectively captures interactions across diverse modalities remains a key challenge in multimodal EHR processing.
Beyond multimodal learning, recent research has explored generative approaches to healthcare modeling. Conventional methods have primarily relied on discriminative models for disease risk assessment and diagnosis Choi et al. (2016); Niu et al. (2021b); Qiao et al. (2019). However, generative models are increasingly being adopted, as demonstrated by Clinical CoT Kwon et al. (2024), applying LLMs for disease diagnosis generation. Reinforcement learning from human feedback (RLHF) Ouyang et al. (2022) and Chain-of-Thought (CoT) prompting Wei et al. (2022) have further enhanced medical reasoning capabilities in models such as GatorTron Yang et al. (2022), MedPalm Singhal et al. (2023), and GPT4-Med Nori et al. (2023). While these models excel in medical question-answering, they remain limited in real-world direct disease diagnosis and multimodal EHR processing. EHR-KnowGen Niu et al. (2024) reframes disease diagnosis as a text-to-text generation problem but overlooks the crucial temporal details embedded in time series lab tests, underscoring the need for more effective and dedicated multimodal fusion strategies.
Figure 2: The ProMedTS model comprises three modules: the Time Series Prompt Embedding (TSPE) module, the Multimodal Textual Information Fusion (MTIF) module, and the Self-supervised Learning (SSL) module. The MTIF module utilizes Clinical-BERT to encode medical notes $\bm{M}$, lab test data $\bm{X}$, and anomaly descriptions $\bm{C}$ to generate time series prompt embeddings $\bm{\mathcal{T}}$.
3 Methodology
In this section, we present the ProMedTS framework for unifying heterogeneous EHR data through prompt-guided learning. We begin by defining the problem and describing the model inputs, then provide a high-level overview of the architecture. In subsequent sections, we detail each module and discuss how these components are applied to downstream tasks such as disease diagnosis.
3.1 Problem Definition
We introduce ProMedTS, aiming to reduce discrepancies between language and time series data in EHRs. Specifically, it leverages anomaly captions and generates time series prompt embeddings to unify both modalities in a shared latent space. The inputs to ProMedTS, denoted by $\{\bm{M},\bm{X}\}$, include medical notes $\bm{M}\in\mathbb{R}^{B\times N_{m}}$ (where $B$ is the batch size and $N_{m}$ is the number of tokens) and numeric lab test data $\bm{X}\in\mathbb{R}^{B\times L\times N_{x}}$ (where $L$ is the sequence length and $N_{x}$ is the number of lab test variants). Additionally, a lightweight anomaly detection Vinutha et al. (2018) is employed to generate textual descriptions of anomalies $\bm{C}\in\mathbb{R}^{B\times N_{c}}$ (details in Appendix A.2). ProMedTS also uses learnable time series query embeddings $\bm{P}\in\mathbb{R}^{B\times N_{p}\times D}$, which are transformed into time series prompt embeddings $\bm{\mathcal{T}}\in\mathbb{R}^{B\times N_{p}\times D}$, where $N_{p}$ is the query length and $D$ is the hidden dimension.
3.2 Model Overview
Figure 2 illustrates the overview of ProMedTS, which comprises three main modules. Three modules share the same Clinical-BERTAlsentzer et al. (2019) structured model and are extended to support cross-attention, self-attention, and prompt generation. The Time Series Prompt Embedding (TSPE) module applies a cross-attention mechanism to convert raw lab test data into prompt embeddings, preserving key temporal features. The Multimodal Textual Information Fusion (MTIF) module encodes and merges medical notes with anomaly captions in a unified latent space, facilitating the extraction of complementary semantic information. Finally, the Self-supervised Learning (SSL) module employs tailored loss functions to bridge the modality gap and maintain fine-grained temporal details in the learned representations. These modules work in tandem to achieve robust alignment and fusion of heterogeneous EHRs, and the following sections provide in-depth explanations of each component and their applications.
3.3 Time Series Prompt Embedding
The objective of the TSPE module is to extract and encapsulate the inherent information from time series lab test data into a time series prompt embedding. Let $\{\bm{X},\bm{P}\}$ represent the module inputs. The numeric lab test data $\bm{X}$ is first processed by a time series encoder (TSE) using PatchTST Nie et al. (2022). In parallel, the learnable query embeddings $\bm{P}$,initialized using vectors extracted from the Clinical-BERT word embedding layer, serve as query tokens in the cross-attention mechanism, guiding the selection of relevant temporal features by attending to time series lab test results encoded by TSE. To generate the final prompt embedding $\bm{\mathcal{T}}$, we extend the multi-head self-attention encoder of Clinical-BERT to support a multi-head cross-attention mechanism, following a strategy similar to that adopted in Li et al. (2023). We designate $\bm{X}$ as both key and value while $\bm{P}$ serves as the query:
$\bm{\mathcal{T}}=\text{Clinical-BERT}\bigl{(}\bm{P},\text{TSE}(\bm{X}),\text{TSE}(\bm{X})\bigr{)}.$ (1)
This design ensures that the rich temporal patterns in $\bm{X}$ are captured within $\bm{\mathcal{T}}$, enabling subsequent modules to leverage these features effectively.
3.4 Multimodal Textual Information Fusion
The MTIF module is designed to fuse medical notes and anomaly descriptions effectively. We use the anomaly captioning method to generate anomaly descriptions, as illustrated in Figure 2. The inputs to the MTIF module are medical notes $\bm{M}$ and lab test anomaly descriptions $\bm{C}$, which are encoded separately by Clinical-BERT via the multi-head self-attention mechanism:
$\bm{E}_{m}=\text{Clinical-BERT}(\bm{M},\bm{M},\bm{M}),$ (2)
$\bm{E}_{c}=\text{Clinical-BERT}(\bm{C},\bm{C},\bm{C}),$
where $\bm{E}_{m}\in\mathbb{R}^{B\times N_{m}\times D}$ and $\bm{E}_{c}\in\mathbb{R}^{B\times N_{c}\times D}$. The repeated inputs indicate that the key, query, and value matrices are identical for the self-attention mechanism. This structure enables the model to encode each type of textual information independently while capturing the inherent characteristics and context of each input. The combined textual representation is then derived from these encoded inputs:
$\bm{E}_{f}=AVG([\bm{E}_{m}\oplus\bm{E}_{c}]),$ (3)
where $\bm{E}_{f}\in\mathbb{R}^{B\times D}$, with $\oplus$ indicating concatenation, and $AVG$ representing average pooling.
3.5 Self-Supervised Learning
This module addresses the modality gap between textual and time series EHR data using three specialized loss functions. By simultaneously aligning cross-modal representations and preserving fine-grained temporal details, the model learns to capture both semantic and temporal nuances.
3.5.1 Cross-Modal Contrastive Alignment
To promote cross-modal alignment, we design a contrastive loss that brings language and time series embeddings closer when they originate from the same patient and pushes them apart otherwise. We first compute similarity matrices by multiplying the fused text representation $\bm{E}_{f}$ with the time series prompt embeddings $\bm{\mathcal{T}}$:
$\bm{g}_{(\bm{E}_{f},\bm{X})}$ $=\max\Bigl{(}\bigl{[}\bm{E}_{f}\,\bm{\mathcal{T}}^{(1)T},\ldots,\bm{E}_{f}\,\bm{\mathcal{T}}^{(N_{p})T}\bigr{]}\Bigr{)},$ (4)
$\bm{g}_{(\bm{X},\bm{E}_{f})}$ $=\max\Bigl{(}\bigl{[}\bm{\mathcal{T}}^{(1)}\,\bm{E}_{f}^{T},\ldots,\bm{\mathcal{T}}^{(N_{p})}\,\bm{E}_{f}^{T}\bigr{]}\Bigr{)},$
where the $\max$ operator performs max-pooling across $N_{p}$ dimensions, yielding $\bm{g}_{(\bm{E}_{f},\bm{X})}$ and $\bm{g}_{(\bm{X},\bm{E}_{f})}\in\mathbb{R}^{B\times B}$. Note that $\bm{g}_{(\bm{E}_{f},\bm{X})}$ measures text-to-time series similarity (by fixing $\bm{E}_{f}$ and iterating over $\bm{\mathcal{T}}$), while $\bm{g}_{(\bm{X},\bm{E}_{f})}$ captures time-series-to-text similarity (by fixing $\bm{\mathcal{T}}$ and iterating over $\bm{E}_{f}$). We then apply the SoftMax function to generate two distinct sets of logits:
$\hat{\bm{y}}_{c}^{f2x}$ $=\text{SoftMax}\bigl{(}\bm{g}_{(\bm{E}_{f},\bm{X})}\bigr{)},$ (5)
$\hat{\bm{y}}_{c}^{x2f}$ $=\text{SoftMax}\bigl{(}\bm{g}_{(\bm{X},\bm{E}_{f})}\bigr{)}.$
Let $\bm{y}_{c}^{f2x}$ and $\bm{y}_{c}^{x2f}$denote the ground truth labels indicating whether the pairs correspond to the same patient in a training batch (1 if matched, 0 otherwise). We use cross-entropy $H(\cdot)$ to define the contrastive loss:
$\mathcal{L}_{contrast}$ $=\;\tfrac{1}{2}\,\mathbb{E}\Bigl{[}H\bigl{(}\bm{y}_{c}^{f2x},\,\hat{\bm{y}}_{c}^{f2x}\bigr{)}\;$ (6)
$+\;H\bigl{(}\bm{y}_{c}^{x2f},\,\hat{\bm{y}}_{c}^{x2f}\bigr{)}\Bigr{]}.$
3.5.2 Intra-Modal Matching
To further capture intra-modality consistency, we align lab tests with corresponding anomaly descriptions. This alignment is modeled as a binary classification task, distinguishing matched from unmatched pairs of lab tests and anomaly captions. Following Li et al. (2021), we employ a negative mining strategy to generate labels $\bm{y}_{m}$ by selecting the most similar pairs in a training batch as negative samples, where the top 1-ranked pair is labeled as 1 and the others as 0, based on the similarity computed in Equation 4. We employ Clinical-BERT’s cross-attention, where the concatenation of $\bm{C}$ and $\bm{P}$ serves as the query, and the encoded time series $\bm{X}$ is used as both key and value. A Multilayer Perceptron (MLP) classifier with softmax activation, denoted $f_{match}$, predicts the probability $\hat{\bm{y}}_{m}$:
$\hat{\bm{y}}_{m}=f_{match}\Bigl{(}\text{Clinical-BERT}\bigl{(}f_{\mathcal{W}}(\bm{C})\oplus\bm{P},$ (7)
$\text{TSE}(\bm{X}),\text{TSE}(\bm{X})\bigr{)}\Bigr{)},$
where $f_{\mathcal{W}}$ is the word embedding layer in Clinical-BERT. We define the matching loss as:
$\mathcal{L}_{match}\;=\;\mathbb{E}\bigl{[}H\bigl{(}\bm{y}_{m},\,\hat{\bm{y}}_{m}\bigr{)}\bigr{]},$ (8)
where $\bm{y}_{m}$ is the one-hot ground truth label.
3.5.3 Anomaly Description Reconstruction
To ensure the time series prompt embeddings encode both coarse anomaly descriptions and fine-grained temporal details, we reconstruct anomaly captions from the learned embeddings. This step helps unify language tokens and time series representations in a shared space. Specifically, we use Clinical-BERT with a language model head $f_{head}$, setting $\bm{E}_{c}$ as the query and $\bm{\mathcal{T}}$ as key and value:
$\mathcal{L}_{gen}=\mathbb{E}\Bigl{[}H\bigl{(}\bm{C},f_{head}\bigl{(}\text{Clinical-BERT}(\bm{E}_{c},\bm{\mathcal{T}},\bm{\mathcal{T}})\bigr{)}\bigr{)}\Bigr{]}.$ (9)
This objective encourages the model to generate accurate textual descriptions, thereby reinforcing alignment between time series prompts and language tokens.
Overall Loss: We combine these objectives into a single training loss:
$\mathcal{L}_{total}=\alpha\,\mathcal{L}_{contrast}\;+\;\beta\,\mathcal{L}_{match}\;+\;\gamma\,\mathcal{L}_{gen},$ (10)
where $\alpha$, $\beta$, and $\gamma$ are hyperparameters balancing the three losses (see Appendix A.6). Our training algorithm aims to minimize $\mathcal{L}_{total}$ across all samples (details in Appendix A.1).
3.6 LLM-based Disease Diagnosis with ProMedTS
To illustrate the practical effectiveness of ProMedTS in unifying textual and time series data, we employ a pre-trained, frozen LLM model for disease diagnosis. As depicted in Figure 3, ProMedTS first converts numeric lab test results into time series prompt embeddings, which are then aligned via a fully connected layer to match the LLM’s input dimensions. These embeddings serve as prefix soft prompts, concatenated with the medical notes so that the model can ingest structured signals from time series alongside unstructured clinical text. By bridging language and time series modalities, the LLM can process both inputs concurrently, leveraging complementary information for enhanced diagnostic accuracy.
Figure 3: ProMedTS for empowering LLMs to in disease diagnosis.
Models Size Type Modality Micro Macro
CLS GEN Lab Note Precision Recall F1 Precision Recall F1
MIMIC-III
GRU 7.9M 46.41(3.48) 21.88(3.59) 29.43(1.89) 30.47(4.23) 13.00(1.14) 14.59(1.48)
PatchTST 19.2M 32.64(3.59) 42.72(5.01) 36.02(1.09) 26.86(3.51) 29.71(4.78) 19.25(3.50)
TimeLLM 78M 37.43(1.17) 54.93(6.56) 36.59(1.17) 10.18(2.30) $35.21_{(6.47)}$ 15.16(2.17)
CAML 36.1M 69.04(0.18) 55.87(2.72) 61.54(0.30) 65.08(2.56) 50.12(3.05) 54.42(0.94)
DIPOLE 39M 64.38(0.89) 57.94(1.15) 60.98(0.27) 61.63(1.03) 53.02(1.18) 55.68(0.49)
PROMPTEHR 75.2M 59.29(0.97) 65.53(0.69) 62.24(0.23) 57.44(0.97) 62.87(0.61) 59.10(0.24)
LLaMA 7B 61.32(2.10) 65.88(1.54) 63.44(0.37) 60.98(1.54) 61.59(1.36) 60.34(0.38)
LDAM 41.3M 68.00(1.23) 57.12(0.47) 62.18(0.40) 67.38(0.35) 51.50(0.95) 57.44(0.60)
FROZEN 265M 61.09(1.81) 64.07(1.58) 62.51(0.34) 59.96(1.55) 59.99(1.66) 59.15(0.30)
EHR-KnowGen 76.9M 60.01(0.29) 65.51(0.18) 62.62(0.06) 58.34(0.38) 61.81(0.28) 59.44(0.06)
ProMedTS 267.5M 61.32(0.54) 66.65(0.51) 63.67(0.08) 60.35(0.61) 61.62(0.71) 60.42(0.18)
ProMedTS* 1B 60.62(0.22) 67.83(0.18) 64.02(0.11) 59.43(0.37) 63.65(0.54) 60.78(0.13)
MIMIC-IV
GRU 7.9M 56.23(1.13) 25.77(1.58) 35.21(1.36) 38.37(1.90) 16.97(1.22) 20.65(1.32)
PatchTST 19.2M 27.26(0.03) 57.42(0.41) 36.97(0.10) 20.59(2.76) 43.72(0.25) 21.78(2.83)
TimeLLM 78M 30.30(1.78) 60.46(1.98) 40.31(1.20) 24.61(2.21) 47.26(2.37) 25.56(1.60)
CAML 36.1M 72.82(0.54) 59.48(0.82) 65.40(0.36) 67.25(0.99) 50.73(1.49) 54.71(1.42)
DIPOLE 39M 72.39(0.51) 61.38(0.83) 66.43(0.33) 70.45(0.37) 55.65(0.79) 60.37(0.62)
PROMPTEHR 75.2M 65.24(0.68) 70.31(0.56) 68.02(0.17) 63.53(0.47) 67.02(0.65) 65.01(0.28)
LLaMA 7B 68.44(1.03) 69.49(0.70) 69.13(0.31) 67.45(0.93) 66.18(1.12) 66.15(0.56)
LDAM 41.3M 72.01(0.85) 62.74(0.62) 66.91(0.20) 69.77(0.18) 56.72(0.69) 60.77(0.48)
FROZEN 265M 67.81(0.78) 69.08(0.94) 68.42(0.08) 66.27(1.00) 65.21(0.97) 65.30(0.05)
EHR-KnowGen 76.9M 65.80(0.64) 70.85(0.45) 68.16(0.11) 63.82(0.53) 67.24(0.55) 65.11(0.13)
ProMedTS 267.5M 71.63(0.46) 67.81(0.85) 69.69(0.18) 70.12(0.47) 63.58(0.79) 66.21(0.17)
ProMedTS* 1B 71.12(0.31) 69.33(0.42) 70.21(0.05) 70.97(0.42) 65.51(0.64) 67.56(0.09)
Table 1: The performance of comparative methods in the disease diagnosis tasks on MIMIC-III and MIMIC-IV. Please note CLS - classification model, GEN -generative model, Lab - lab test result, and Note - medical notes.
4 Experiments
4.1 Datasets and Preprocessing
The MIMIC-III dataset Johnson et al. (2016) is a publicly available EHR dataset containing de-identified patients who were admitted to ICUs between 2001 and 2012. It includes medical discharge summaries, lab test results, chest x-ray images and more. Our analysis focuses on EHR data from approximately 27,000 patients including complete medical discharge summaries and lab test results. The MIMIC-IV dataset Johnson et al. (2023) comprises EHR data from 2008 to 2019. We utilize approximately 29,000 EHR records from MIMIC-IV, which include complete medical discharge summaries and lab test results. Our study targets 25 disease phenotypes as defined in the MIMIC-III benchmark Harutyunyan et al. (2019a).
Data Pre-processing.
For medical notes, we extract the brief course from medical discharge summaries, removing numbers, noise, and stopwords. Numerical lab test results are converted into time series data using the benchmark tools Harutyunyan et al. (2019b), with missing values filled using the nearest available numbers. Time series anomaly descriptions are used the method defined in Appendix A.2. Data splitting follows the guidelines in Harutyunyan et al. (2019b), using a 4:1 ratio for training and testing.
4.2 Baseline Methods
We benchmark our approach against a range of methods: GRU Cho et al. (2014), PatchTST Nie et al. (2022), TimeLLM Jin et al. (2023), CAML Mullenbach et al. (2018), DIPOLE Ma et al. (2017), PROMPTEHR Wang and Sun (2022), LLaMA–7B Touvron et al. (2023), LDAM Niu et al. (2021a), FROZEN Tsimpoukelli et al. (2021), and EHR-KnowGen Niu et al. (2024). Detailed configurations of these baselines are provided in Appendix A.3. For the disease diagnosis task, we adopt two scales of Flan-T5 Chung et al. (2024) as the inference LLM to validate our model’s effectiveness in understanding multimodal EHRs, primarily driven by resource considerations and ease of experimentation. The Flan-T5-Small-based model is denoted as PromMedTS, while the Flan-T5-Large-based model is denoted as ProMedTS*. In principle, any sufficiently LLM could be substituted to potentially achieve even stronger results.
To ensure a fair comparison, all baselines also employ Flan-T5 as their backbone. Reported results are averaged over five runs with different random seeds. The statistical significance determined at p < 0.05 by t-test. Implementation details for every model are described in Appendix A.4, and training instructions appear in Appendix A.5. Our code is publicly available at https://anonymous.4open.science/r/ProMedTS-V1-5F51/.
Models Micro Macro
Precision Recall F1 Precision Recall F1
MIMIC-III
ProMedTS 61.32(0.54) 66.65(0.51) 63.67(0.08) 60.35(0.61) 61.62(0.71) 60.42(0.18)
w/o LAB 58.91(0.83) 66.59(0.57) 62.34(0.26) 57.32(0.88) 62.56(0.61) 59.05(0.22)
w/o ANOMALY 60.09(0.32) 65.03(0.98) 62.44(0.22) 59.13(0.43) 60.46(1.15) 59.11(0.25)
MIMIC-IV
ProMedTS 71.63(0.46) 67.81(0.85) 69.69(0.18) 70.12(0.47) 63.58(0.79) 66.21(0.17)
w/o LAB 67.16(0.55) 69.42(0.59) 68.22(0.31) 65.74(0.62) 64.69(0.48) 64.33(0.18)
w/o ANOMALY 70.94(0.37) 66.44(1.37) 68.47(0.12) 68.95(0.79) 62.45(0.96) 65.13(0.12)
Table 2: Ablation studies on different modality input and alignment designs for disease diagnosis.
Models Micro Macro
Precision Recall F1 Precision Recall F1
MIMIC-III
ProMedTS 61.32(0.54) 66.65(0.51) 63.67(0.08) 60.35(0.61) 61.62(0.71) 60.42(0.18)
w/o CONTRAST 60.24(0.25) 66.00(0.39) 62.99(0.07) 59.92(0.60) 61.41(0.51) 59.73(0.07)
w/o MATCH 60.12(0.58) 66.14(1.50) 62.96(0.02) 59.70(1.18) 61.37(1.34) 59.65(0.11)
w/o GEN 59.95(0.38) 66.15(0.27) 62.89(0.19) 59.57(0.55) 61.32(0.30) 59.61(0.20)
MIMIC-IV
ProMedTS 71.63(0.46) 67.81(0.85) 69.69(0.18) 70.12(0.47) 63.58(0.79) 66.21(0.17)
w/o CONTRAST 70.19(0.25) 66.22(0.39) 68.61(0.09) 69.05(0.24) 62.40(0.35 65.21(0.12)
w/o MATCH 70.79(0.34) 66.49(0.38) 68.67(0.15) 68.91(0.73) 62.25(0.48) 65.47(0.15)
w/o GEN 71.30(0.29) 65.79(0.57) 68.44(0.13) 69.14(0.48) 62.05(0.54) 65.03(0.13)
Table 3: Ablation studies on the effectiveness of different loss functions of our model for disease diagnosis.
4.3 Disease Diagnosis Performance
Table 1 shows that ProMedTS achieves the highest overall performance, particularly in F1 scores on MIMIC-IV. In addition, replacing the LLM with a larger model improves F1 scores on both datasets, indicating our model’s scalability and robustness across different LLMs. Furthermore, TimeLLM performs strongly with lab test, highlighting the value of time-series inputs for LLMs in disease diagnosis. Text-based methods (e.g., LLaMA) generally outperform time-series approaches, suggesting that medical notes capture richer disease-related information. Multimodal models (LDAM, FROZEN, EHR-KnowGen) exceed single-modality baselines (TimeLLM, CAML, DIPOLE, PROMPTEHR), confirming the benefits of integrating text and time series. Generative approaches (TimeLLM, LLaMA, FROZEN, EHR-KnowGen) also outperform classification-based methods. Although LLaMA performs well, its higher variance and parameter requirements reduce practicality. Notably, the Flan-T5-based ProMedTS and ProMedTS* surpass all baselines in F1 on both datasets, highlighting its efficiency and effectiveness.
4.4 Ablation Studies
4.4.1 Effect of Modality Alignment in ProMedTS
This section presents ablation studies to evaluate each module in ProMedTS. ProMedTS w/o LAB excludes lab test, removing modality alignment with anomaly descriptions and medical notes. ProMedTS w/o ANOMALY removes alignment with anomaly descriptions while keeping alignment between lab test and medical notes to assess the impact of self-supervision. Table 2 summarizes the results, showing that ProMedTS w/o LAB suffers a significant drop in F1 scores, highlighting the importance of lab test. ProMedTS w/o ANOMALY also shows reduced performance, highlighting the challenges of aligning modalities from discrete and continuous encoding spaces and the adverse effects of misalignment on multimodal understanding.
4.4.2 Impact of Self-Supervised Loss Functions
Table 3 summarizes an ablation study on the loss functions in ProMedTS. Both ProMedTS w/o CONTRAST and ProMedTS w/o MATCH show slight declines in F1 scores, emphasizing the importance of $\mathcal{L}_{contrast}$ for aligning and unifying time series and textual inputs within a shared latent space. The results also underscore the role of $\mathcal{L}_{match}$ in intra-modal alignment, ensuring the distinctiveness of time series data by aligning lab test with time series prompt embeddings. Notably, ProMedTS w/o GEN exhibits a significant drop in F1 scores, highlighting the critical role of $\mathcal{L}_{gen}$ in refining prompt embeddings and integrating temporal information from time series data and anomaly descriptions.
Figure 4: The model parameters and computation time of all baselines.
4.5 Model Efficiency and Complexity
Figure 4 illustrates the parameter counts and computation times of baseline models on the two datasets. Our model, ProMedTS, matches the parameter counts and computation times of multimodal baselines such as LDAM and FROZEN, while using 25× fewer parameters and requiring one-third less training time than LLaMA, all while achieving superior diagnostic performance, highlighting its efficiency and effectiveness in language-time series multimodal alignment and fusion.
      $N_{p}$       Micro F1       Macro F1
       MIMIC-III
      12       63.09(0.06)       59.69(0.12)
      24       63.67(0.08)       60.42(0.18)
      36       63.32(0.09)       59.96(0.15)
       MIMIC-IV
      12       68.98(0.15)       65.43(0.18)
      24       69.69(0.18)       66.21(0.17)
      36       69.41(0.19)       65.91(0.20)
Table 4: Sensitivity analysis on different length of time series prompt embedding.
4.6 Sensitivity Analysis of Time Series Prompt Length
We performed a sensitivity analysis to examine the impact of the time series prompt embedding length ($N_{p}$) on the performance of ProMedTS in disease diagnosis. Table 4 shows the F1 scores for embedding lengths of 12, 24, and 36. Slight fluctuations are observed in both micro and macro F1 scores across the MIMIC-III and MIMIC-IV datasets. The optimal embedding length is 24 for both datasets, consistent with the configuration used in our experiments.
Lab Test Input Micro F1 Macro F1
MIMIC-III
Numerical Values 32.21(1.33) 23.53(1.21)
Anomaly captions 35.19(0.92) 24.75(0.76)
Time series prompts 36.11(1.14) 25.47(1.02)
MIMIC-IV
Numerical Values 37.75(1.46) 26.10(1.09)
Anomaly captions 39.56(1.14) 27.22(0.77)
Time series prompts 40.14(1.05) 28.43(0.91)
Table 5: Micro and Macro F1 Scores Across Various Lab Test Input Types on Flan-T5 for Disease Diagnosis
4.7 Evaluating the Role of Anomaly Descriptions
To highlight the advantages of using lab test anomaly captions over raw numerical time series values in LLMs, we evaluate Flan-T5-small with both input types. Table 5 presents the evaluation results on the MIMIC-III and MIMIC-IV datasets for disease diagnosis. The results show that Flan-T5 achieves over a 2% improvement in Micro F1 score when using anomaly captions, demonstrating that LLMs interpret anomaly captions more effectively than raw numerical values in time series lab test data. Additionally, the inclusion of time series prompts underscores the effectiveness of our model, ProMedTS, in capturing both fine-grained and coarse-grained temporal information from lab test results for disease diagnosis.
5 Conclusion and Future Work
In this paper, we introduce ProMedTS, a lightweight and effective modality fusion framework that leverages self-supervised prompt learning for multimodal EHR integration. By bridging the modality gap between medical notes and lab test results, ProMedTS enables LLMs to process structured and unstructured medical data more effectively. Its three key modules and self-supervised loss functions advance language–time series integration in healthcare, providing a scalable and adaptable approach for real-world clinical applications. Evaluation on two EHR datasets demonstrates that ProMedTS significantly outperforms existing models in disease diagnosis, underscoring its potential to enhance clinical decision-making and improve patient care. In future work, we plan to extend our approach to larger and more diverse datasets, explore additional LLM architectures, and investigate further improvements in modality alignment techniques.
Limitations
While this study focuses on modality alignments and their application in downstream tasks, enhancing the explainability of disease diagnosis remains an area for future work, where we plan to incorporate the Chain-of-Thought rationale Wei et al. (2022). Additionally, computational constraints required the use of a relatively compact LLM, limiting the amount of clinical text processed at once, which may impact the model’s ability to leverage full medical histories. Expanding to more capable models will help address this challenge. Furthermore, our study primarily targets higher-level disease phenotypes in the International Classification of Diseases (ICD) codes Slee (1978), which could be expended to more downstream tasks. Future work will explore larger models, such as LLaMA Touvron et al. (2023) and Mistral Jiang et al. (2023), to improve diagnostic granularity and broaden coverage.
Ethics Statement
Data Privacy:
While the datasets utilized in our research, such as MIMIC-III and MIMIC-IV, are publicly accessible and feature de-identified patient data, accessing these datasets still requires passing the CITI examination and applying for the data through PhysioNet.
Appendix A Appendix
A.1 Lab Test Anomaly Caption
Time series anomaly descriptions are generated using the IQR method Vinutha et al. (2018) to identify anomalies, capturing their timing and polarity (above or below standard values) and describing them with handcrafted templates. To caption the lab test anomaly into textual format, we design several text templates to describe the lab test anomalies. All templates are illustrated in Table 6.
A.2 Algorithm
The training procedure to optimize ProMedTS by minimizing the loss defined in Equation (10) is shown in Algorithm 1.
Algorithm 1 The ProMedTS Model
1:  Input: Given lab test $\bm{X}$ and medical note $\bm{M}$ denote the EHRs input. $\bm{P}$ consists of a set of learnable prompt embeddings.
2:  while not converge do
3:     for mini-batch $B$ do
4:        Obtain the time series anomaly caption $\bm{\mathcal{T}}$ using equation (1).
5:        Obtain multimodal textual embedding $\bm{E}_{f}$ using equations (2) and (3).
6:        Calculate the contrastive loss $\mathcal{L}_{contrast}$ between lab test, anomalies, and medical notes using equations (4), (5), and (6).
7:        Calculate the matching loss $\mathcal{L}_{match}$ between lab test and anomalies using equations (7) and (8).
8:        Calculate the generation loss $\mathcal{L}_{gen}$ between lab test and anomalies using equation (9).
9:     end for
10:     Update parameters by minimizing the total loss $\mathcal{L}_{total}$ defined in Equation (10) by using the AdamW optimizer Loshchilov and Hutter (2018) for patients in each batch.
11:  end while
A.3 Baseline Models
If lab test value is not an abnormal value
{Lab features} is normal all the time.
If the lab test value is an abnormal value higher than the standard.
{Lab features} is higher than normal {number of times} times.
If the lab test value is an abnormal value lower than the standard.
{Lab features} is lower than normal {number of times} times.
If the lab test value is an abnormal that include both higher and lower than the standard value.
{Lab features} is higher than normal {number of times} times and lower than normal {number of
times} times.
Table 6: Lab test anomaly caption template.
  • GRU: The Gated Recurrent Unit (GRU) Cho et al. (2014), a variant of recurrent neural networks (RNNs), employs two gates to capture both long-term and short-term temporal features effectively.
  • PatchTST: PatchTST Nie et al. (2022) is a transformer-based time series encoder designed for long-term forecasting. It segments time series into subseries-level patches, treating each as a token within the transformer architecture.
  • TimeLLM: TimeLLM Jin et al. (2023) is an LLM-based time series prediction model that reprograms input time series into text prototypes before processing them with a frozen LLM. It achieves state-of-the-art performance in mainstream forecasting tasks, particularly in few-shot and zero-shot scenarios.
  • CAML: The Convolutional Attention for Multi-Label classification (CAML) Mullenbach et al. (2018) is a classical model for classifying medical notes, incorporating a cross-attention mechanism and label embeddings to enhance interpretability. For a fair comparison with more recent language models, its original embedding layer is replaced with one from T5.
  • DIPOLE: DIPOLE Ma et al. (2017) is a classic disease prediction model that utilizes two Bi-directional RNNs. It incorporates an attention mechanism to integrate information from both past and future hospital visits. For a fair comparison with more recent language models, its original embedding layer has been replaced with one from T5.
    Diagnose disease from the following medical notes and lab test:
    Medical Notes: ms woman signficant pmh atrial fibrillation lung cancer resection congestive heart
    hypertension worsening diarrhea dysuria hypotensive admitted unit presumed active issues hypotensive
    pronounced leukocytosis multiple potential sources ct scan peritoneal fluids free fluid started surgery …
    Lab Test: $<prefix_{1}>,<prefix_{2}>,...,<prefix_{n}>$
    Diagnosis:
    Acute and unspecified renal failure, Fluid and electrolyte disorders, Septicemia (except in labor), Shock,
    Chronic obstructive pulmonary disease and bronchiectasis, Disorders of lipid metabolism,
    Cardiac dysrhythmias, Congestive heart failure; nonhypertensive, Diabetes mellitus with complications,
    Other liver diseases.
    Table 7: Training Instruction Template
  • PROMPTEHR: PROMPTEHR Wang and Sun (2022) introduces a novel approach in generative models for electronic health records (EHRs), implementing conditional prompt learning. In this study, the model is specifically geared towards disease diagnosis.
  • LLaMA: LLaMA-7B Touvron et al. (2023), one of the leading large language models, is enhanced by Reinforcement Learning with Human Feedback (RLHF) and instructive tuning. It is fine-tuned for disease diagnosis in this study, demonstrating its versatility in various NLP tasks.
  • LDAM: LDAM Niu et al. (2021a) leverages multimodal inputs, combining laboratory testing results and medical notes for disease risk prediction. It utilizes label embedding to effectively integrate these two modalities.
  • FROZEN: FROZEN Tsimpoukelli et al. (2021) represents the cutting-edge multimodal vision-language models for few-shot learning. In our study, it is adapted to the disease diagnosis task using inputs from lab test results and medical notes.
  • EHR-KnowGen: EHR-KnowGen Niu et al. (2024), touted as the state-of-the-art in EHR multimodal learning models, focuses on disease diagnosis generation. For this study, external domain knowledge is excluded to ensure a fair comparison.
Figure 5: Sensitivity analysis of varying ratios in loss function components.
A.4 Implementation Details
In experiments, we utilized PyTorch framework version 2.0.1, operating on a CUDA 11.7 environment. We employed the AdamW optimizer with a starting learning rate of $1e^{-5}$ and a weight decay parameter of 0.05. Additionally, we implemented a warm-up strategy covering 10% of the training duration. Our experiments were conducted on high-performance NVIDIA Tesla V100 GPUs. Within the ProMedTS model, we used 24 time series prompt embeddings, each with a dimensionality of 768. The model’s hidden layer size was maintained at 768 for modality alignment and adjusted to 512 for downstream tasks. To standardize the time series data input, we padded all lab test results to a uniform length of 1000 time steps, allowing us to divide the data into 125 patches, with each patch containing 8 time steps. The Flan-T5 is fine-tuned on two MIMIC datasets Johnson et al. (2016, 2023) and then frozen for downstream tasks.
A.5 Training Instruction Template
Table 7 illustrates the training instruction template for our model ProMedTS for disease diagnosis on MIMIC-III and MIMIC-IV datasets.
A.6 Sensitivity Analysis of Varying Ratios in Loss Function Components
To examine the impact of different combinations of the three loss functions, $\mathcal{L}_{contrast}$, $\mathcal{L}_{match}$, and $\mathcal{L}_{gen}$, on the downstream performance, we perform a sensitivity analysis using three sets of loss ratios: 1:1:1, 1:2:2, and 1:2:1 on MIMIC-III and MIMIC-IV datasets. Since the value of $\mathcal{L}_{contrast}$ is typically larger than those of $\mathcal{L}_{match}$ and $\mathcal{L}_{gen}$, we assign greater weights to $\mathcal{L}_{match}$ and $\mathcal{L}_{gen}$. Figure 5 presents the results, where lines indicate the variation in the sum of the three loss functions on the testing dataset and bars represent the Micro and Macro F1 scores. The figure reveals that varying the weight ratios of the three loss functions has minimal impact on model convergence and the performance of downstream disease diagnosis tasks.

### v2

@@
-Unlocking Multimodal Integration in EHRs: A Prompt Learning Framework for Language and Time Series Fusion
+ProMedTS: A Self-Supervised, Prompt-Guided Multimodal Approach for Integrating Medical Text and Time Series
 Abstract
 Large language models (LLMs) have shown remarkable performance in vision-language tasks, but their application in the medical field remains underexplored, particularly for integrating structured time series data with unstructured clinical notes.
-In clinical practice, dynamic time series data such as lab test results capture critical temporal patterns, while clinical notes provide rich semantic context.
+In clinical practice, dynamic time series data, such as lab test results, capture critical temporal patterns, while clinical notes provide rich semantic context.
 Merging these modalities is challenging due to the inherent differences between continuous signals and discrete text.
 To bridge this gap, we introduce ProMedTS, a novel self-supervised multimodal framework that employs prompt-guided learning to unify these heterogeneous data types.
-Our approach leverages lightweight anomaly detection to generate anomaly captions that serve as prompts, guiding the encoding of raw time series data into informative embeddings.
-These embeddings are aligned with textual representations in a shared latent space, preserving fine-grained temporal nuances alongside semantic insights.
+Our approach leverages lightweight anomaly detection to generate anomaly captions that serve as prompts, guiding the encoding of raw time series data into informative prompt embeddings.
+These prompt embeddings are aligned with textual representations in a shared latent space, preserving fine-grained temporal nuances alongside semantic insights.
 Furthermore, our framework incorporates tailored self-supervised objectives to enhance both intra- and inter-modal alignment.
 We evaluate ProMedTS on disease diagnosis tasks using real-world datasets, and the results demonstrate that our method consistently outperforms state-of-the-art approaches.
-Unlocking Multimodal Integration in EHRs: A Prompt Learning Framework for Language and Time Series Fusion
-Shuai Niu1, Jing Ma1, Hongzhan Lin1, Liang Bai2, Zhihua Wang3, Wei Bi5, Yida Xu1, Guo Li4, and Xian Yang6thanks: * This is the corresponding author.
-Hong Kong Baptist University1, Shanxi University2, Shanghai Institute for Advanced Study of Zhejiang University3, Manchester Metropolitan University4, Tencent AI Lab5, The University of Manchester6 {cssniu,majing, cshzlin}@comp.hkbu.edu.hk, sxbailiang@126.com, zhihua.wang@zju.edu.cn, victoriabi@tencent.com, xuyida@hkbu.edu.hk, l.guo@mmu.ac.uk, xian.yang@manchester.ac.uk
 Figure 1: (a) LLMs struggle to process continuous time series data due to modality gaps with discrete textual representations.
@@
 (2022); Chen et al.
-(2021), integrating continuous time series data with text remains a challenge.
-Unlike text, which is composed of discrete tokens, time series data contain continuous signals with temporal dependencies Jin et al.
-(2023) as illustrated in Figure 1(a).
+(2021), integrating time series data with text remains a challenge.
+Unlike text, which is composed of discrete tokens, time series data contains continuous signals with temporal dependencies as illustrated in Figure 1(a).
 Current multimodal learning approaches, especially contrastive learning methods Radford et al.
 (2021); Li et al.
-(2023), have been effective in aligning vision and language.
-However, they are less suited to bridge the gap between time series and text.
+(2023), have been effective in aligning vision data and text data.
+However, they are less suited to bridge the gap between time series data and text data.
 Time series data require fine-grained temporal representations in a high-dimensional space and are often irregularly sampled, exhibit diverse frequencies, and include missing values Harutyunyan et al.
@@
 In addition, the lack of large-scale paired datasets that link raw time series with textual descriptions further hampers LLMs from incorporating structured information into clinical decision-making Niu et al.
-(2024).
+(2024a).
 Without an effective fusion mechanism, LLMs cannot fully exploit the rich temporal patterns in structured EHR data.
-To address these challenges, we propose ProMedTS, a self-supervised and prompt-guided framework designed to unify medical notes and time series lab test for naturally understood by LLMs.
-As shown in Figure 1(b), instead of feeding raw time series directly into LLMs, our framework uses anomaly descriptions, capturing key patterns in lab test results for helping multimodal fusion between text and time series data.
-These descriptions are generated using lightweight anomaly detection techniques Vinutha et al.
-(2018) and convert continuous signals into human-readable summaries.
-The process involves two steps.
-First, anomaly descriptions establish a direct connection between time series EHRs and medical notes.
-Second, time series prompt embeddings are generated and added as prefix tokens to the LLM input.
-This method integrates structured time series information into the language modeling process without altering the LLM architecture, unifying both modalities within a same encoding space and enhancing clinical decision-making.
-We optimize ProMedTS with three self-supervised learning objectives.
-A contrastive loss maps textual and time series modalities into a shared latent space.
-An anomaly-time series matching loss links lab test with their corresponding anomaly descriptions to reinforce consistency.
-Finally, an anomaly caption generation loss improves the fine-grained alignment between numeric time series lab test and time series prompt embeddings.
+To address these challenges, we propose ProMedTS, a self-supervised and prompt-guided framework designed to unify medical notes and time series lab tests for natural comprehension by LLMs.
+As illustrated in Figure 1(b), rather than only feeding raw time series data directly into LLMs, our framework introduces anomaly descriptions as a modality bridge to facilitate alignment between text and time series data and to aid in generating time series prompt embeddings.
+These descriptions are produced using anomaly detection technology Vinutha et al.
+(2018), which converts continuous signals into human-readable summaries that encompass coarse-grained time series patterns.
+The process consists of two steps.
+First, anomaly descriptions establish a direct connection between time series lab tests and medical notes.
+Second, the generated time series prompt embeddings contain both coarse-grained anomaly information and fine-grained time series variation patterns.
+These embeddings are then appended as prefix tokens to the input of the LLM.
+This approach integrates structured time series information into the language modeling process without modifying the LLM architecture, thereby unifying both modalities within the same encoding space and enhancing clinical decision-making.
+We optimize ProMedTS with three self-supervised learning objectives for learning the time series prompt embeddings.
+A contrastive loss maps textual and time series modalities into a shared latent space to learn the coarse-grained lab test information.
+An anomaly-time series matching loss links numeric lab tests with their corresponding anomaly descriptions to reinforce consistency and learn the fine-grained lab test information.
+Finally, an anomaly caption generation loss enhances the representation of multi-granularity time series information within the time series prompt embeddings.
 Together, these objectives enable LLMs to process both structured and unstructured EHR data more effectively, addressing the gap between language and time series representations in healthcare applications.
@@
 2 Related Work
+2.1 Multimodal Learning in Healthcare
 The increasing diversity of EHR data has led to significant advancements in multimodal learning for healthcare applications.
@@
 (2019) and GLoRIA Huang et al.
-(2021) integrate numerical and image data with text using attention mechanisms.
+(2021) integrate numerical or image data with text using attention mechanisms.
 LDAM Niu et al.
-(2021a) further extends these approaches by leveraging cross-attention with disease labels to fuse features from lab tests and clinical notes.
+(2021) further extends these approaches by leveraging cross-attention with disease labels to fuse features from lab tests and clinical notes.
 EHR-KnowGen Niu et al.
-(2024) transforms structured lab data into text and incorporates external knowledge for improved modality fusion.
+(2024a) transforms structured lab data into text and incorporates external knowledge for improved modality fusion.
 Despite these advancements, achieving a unified latent embedding that effectively captures interactions across diverse modalities remains a key challenge in multimodal EHR processing.
+2.2 Healthcare with LLMs
 Beyond multimodal learning, recent research has explored generative approaches to healthcare modeling.
@@
 EHR-KnowGen Niu et al.
-(2024) reframes disease diagnosis as a text-to-text generation problem but overlooks the crucial temporal details embedded in time series lab tests, underscoring the need for more effective and dedicated multimodal fusion strategies.
+(2024a) reframes disease diagnosis as a text-to-text generation problem but overlooks the crucial temporal details embedded in time series lab tests, underscoring the need for more effective and dedicated multimodal fusion strategies.
 Figure 2: The ProMedTS model comprises three modules: the Time Series Prompt Embedding (TSPE) module, the Multimodal Textual Information Fusion (MTIF) module, and the Self-supervised Learning (SSL) module.
@@
 3.1 Problem Definition
-We introduce ProMedTS, aiming to reduce discrepancies between language and time series data in EHRs.
+We introduce ProMedTS, aiming to reduce discrepancies between language and time series EHRs.
 Specifically, it leverages anomaly captions and generates time series prompt embeddings to unify both modalities in a shared latent space.
@@
 Three modules share the same Clinical-BERTAlsentzer et al.
-(2019) structured model and are extended to support cross-attention, self-attention, and prompt generation.
+(2019) structured model for clinical tokens embedding and are extended to support cross-attention, self-attention, and prompt generation.
 The Time Series Prompt Embedding (TSPE) module applies a cross-attention mechanism to convert raw lab test data into prompt embeddings, preserving key temporal features.
 The Multimodal Textual Information Fusion (MTIF) module encodes and merges medical notes with anomaly captions in a unified latent space, facilitating the extraction of complementary semantic information.
-Finally, the Self-supervised Learning (SSL) module employs tailored loss functions to bridge the modality gap and maintain fine-grained temporal details in the learned representations.
-These modules work in tandem to achieve robust alignment and fusion of heterogeneous EHRs, and the following sections provide in-depth explanations of each component and their applications.
+Finally, the Self-supervised Learning (SSL) module employs tailored loss functions to bridge the modality gap and maintain multi-granularity temporal details in the learned representations.
+These modules work in tandem to achieve robust alignment and fusion of heterogeneous EHRs, and the following sections provide in-depth explanations of each component and its applications.
 3.3 Time Series Prompt Embedding
-The objective of the TSPE module is to extract and encapsulate the inherent information from time series lab test data into a time series prompt embedding.
+The objective of the TSPE module is to extract and encapsulate the inherent fine-grained temporal information from time series lab tests into time series prompt embeddings.
 Let $\{\bm{X},\bm{P}\}$ represent the module inputs.
@@
 (2022).
-In parallel, the learnable query embeddings $\bm{P}$,initialized using vectors extracted from the Clinical-BERT word embedding layer, serve as query tokens in the cross-attention mechanism, guiding the selection of relevant temporal features by attending to time series lab test results encoded by TSE.
+In parallel, the learnable query embeddings $\bm{P}$, initialized using vectors slicing from the Clinical-BERT word embedding layer, serve as query tokens in the cross-attention mechanism, guiding the selection of relevant temporal features by attending to time series lab tests encoded by TSE.
 To generate the final prompt embedding $\bm{\mathcal{T}}$, we extend the multi-head self-attention encoder of Clinical-BERT to support a multi-head cross-attention mechanism, following a strategy similar to that adopted in Li et al.
@@
 This structure enables the model to encode each type of textual information independently while capturing the inherent characteristics and context of each input.
-The combined textual representation is then derived from these encoded inputs:
+The combined textual representation is then derived from $\bm{E}_{m}$ and $\bm{E}_{c}$:
 $\bm{E}_{f}=AVG([\bm{E}_{m}\oplus\bm{E}_{c}]),$ (3)
@@
 This module addresses the modality gap between textual and time series EHR data using three specialized loss functions.
-By simultaneously aligning cross-modal representations and preserving fine-grained temporal details, the model learns to capture both semantic and temporal nuances.
+By simultaneously aligning cross-modal representations and preserving coarse- and fine-grained temporal details, the model learns to capture both semantic and temporal nuances.
 3.5.1 Cross-Modal Contrastive Alignment
-To promote cross-modal alignment, we design a contrastive loss that brings language and time series embeddings closer when they originate from the same patient and pushes them apart otherwise.
+To promote cross-modal alignment and preserve coarse-grained temporal information into time series prompt embeddings, we design a contrastive loss that brings multimodal textual embedding and time series embeddings closer when they originate from the same patient and pushes them apart otherwise.
 We first compute similarity matrices by multiplying the fused text representation $\bm{E}_{f}$ with the time series prompt embeddings $\bm{\mathcal{T}}$:
@@
 Note that $\bm{g}_{(\bm{E}_{f},\bm{X})}$ measures text-to-time series similarity (by fixing $\bm{E}_{f}$ and iterating over $\bm{\mathcal{T}}$), while $\bm{g}_{(\bm{X},\bm{E}_{f})}$ captures time-series-to-text similarity (by fixing $\bm{\mathcal{T}}$ and iterating over $\bm{E}_{f}$).
+This process is the same as that used in vision-language contrastive learning Radford et al.
+(2021); Li et al.
+(2023).
 We then apply the SoftMax function to generate two distinct sets of logits:
@@
 3.5.2 Intra-Modal Matching
-To further capture intra-modality consistency, we align lab tests with corresponding anomaly descriptions.
+To further capture intra-modality consistency and enhance alignment with fine-grained temporal information, we align lab tests with their corresponding anomaly descriptions and time series prompt embeddings.
 This alignment is modeled as a binary classification task, distinguishing matched from unmatched pairs of lab tests and anomaly captions.
@@
 (2021), we employ a negative mining strategy to generate labels $\bm{y}_{m}$ by selecting the most similar pairs in a training batch as negative samples, where the top 1-ranked pair is labeled as 1 and the others as 0, based on the similarity computed in Equation 4.
-We employ Clinical-BERT’s cross-attention, where the concatenation of $\bm{C}$ and $\bm{P}$ serves as the query, and the encoded time series $\bm{X}$ is used as both key and value.
+We employ Clinical-BERT’s cross-attention, where the concatenation of $\bm{C}$ and $\bm{P}$ serves as the query (aiming to align the time series prompt embedding more closely with the textual encoding space during training), and the encoded time series $\bm{X}$ is used as both key and value.
 A Multilayer Perceptron (MLP) classifier with softmax activation, denoted $f_{match}$, predicts the probability $\hat{\bm{y}}_{m}$:
@@
 $\mathcal{L}_{gen}=\mathbb{E}\Bigl{[}H\bigl{(}\bm{C},f_{head}\bigl{(}\text{Clinical-BERT}(\bm{E}_{c},\bm{\mathcal{T}},\bm{\mathcal{T}})\bigr{)}\bigr{)}\Bigr{]}.$ (9)
-This objective encourages the model to generate accurate textual descriptions, thereby reinforcing alignment between time series prompts and language tokens.
+This objective is a standard language model generation loss, computed as cross-entropy between the predicted token distribution and the ground truth tokens, encouraging the model to generate accurate textual descriptions, thereby reinforcing alignment between time series prompts and language tokens.
 Overall Loss: We combine these objectives into a single training loss:
@@
 To illustrate the practical effectiveness of ProMedTS in unifying textual and time series data, we employ a pre-trained, frozen LLM model for disease diagnosis.
-As depicted in Figure 3, ProMedTS first converts numeric lab test results into time series prompt embeddings, which are then aligned via a fully connected layer to match the LLM’s input dimensions.
-These embeddings serve as prefix soft prompts, concatenated with the medical notes so that the model can ingest structured signals from time series alongside unstructured clinical text.
+As depicted in Figure 3, during ProMedTS’s fine-tuning, it first transforms numeric lab test results into time series prompt embeddings, which are subsequently aligned with the LLM’s input dimensions through a tunable fully connected layer.
+These embeddings then serve as prefix soft prompts, concatenated with the medical notes so that the model can ingest structured signals from time series alongside unstructured clinical text.
 By bridging language and time series modalities, the LLM can process both inputs concurrently, leveraging complementary information for enhanced diagnostic accuracy.
@@
 DIPOLE 39M 64.38(0.89) 57.94(1.15) 60.98(0.27) 61.63(1.03) 53.02(1.18) 55.68(0.49)
+Flan-T5 60M 58.12(1.11) 66.23(0.72) 62.03(0.54) 56.56(1.03) 62.47(0.76) 58.87(0.71)
+OPT 125M 42.56(0.95) 69.22(0.87) 52.71(0.71) 47.58(0.97) 65.06(0.85) 52.56(0.68)
+QWEN-2.5 500M 49.77(1.13) 58.24(0.95) 53.21(0.90) 50.32(1.08) 54.03(0.89) 52.99(0.92)
 PROMPTEHR 75.2M 59.29(0.97) 65.53(0.69) 62.24(0.23) 57.44(0.97) 62.87(0.61) 59.10(0.24)
-LLaMA 7B 61.32(2.10) 65.88(1.54) 63.44(0.37) 60.98(1.54) 61.59(1.36) 60.34(0.38)
+LLaMA-1 7B 61.42(.2.08) 65.98(1.53) 63.64(0.41) 61.08(1.54) 61.64(1.27) 60.55(0.44)
 LDAM 41.3M 68.00(1.23) 57.12(0.47) 62.18(0.40) 67.38(0.35) 51.50(0.95) 57.44(0.60)
@@
 EHR-KnowGen 76.9M 60.01(0.29) 65.51(0.18) 62.62(0.06) 58.34(0.38) 61.81(0.28) 59.44(0.06)
-ProMedTS 267.5M 61.32(0.54) 66.65(0.51) 63.67(0.08) 60.35(0.61) 61.62(0.71) 60.42(0.18)
-ProMedTS* 1B 60.62(0.22) 67.83(0.18) 64.02(0.11) 59.43(0.37) 63.65(0.54) 60.78(0.13)
+ProMedTS
+w/ OPT 316M 54.57(0.14) 54.92(0.21) 54.74(0.09) 55.92(0.08) 52.15(0.13) 53.68(0.07)
+w/ QWEN-2.5 685M 55.37(0.15) 58.97(0.18) 57.82(0.13) 56.45(0.14) 56.72(0.16) 55.97(0.13)
+w/ Flan-T5-small 268M 61.32(0.54) 66.65(0.51) 63.67(0.08) 60.35(0.61) 61.62(0.71) 60.42(0.18)
+w/ Flan-T5-large 1B 60.62(0.22) 67.83(0.18) 64.02(0.11) 59.43(0.37) 63.65(0.54) 60.78(0.13)
 MIMIC-IV
@@
 DIPOLE 39M 72.39(0.51) 61.38(0.83) 66.43(0.33) 70.45(0.37) 55.65(0.79) 60.37(0.62)
+Flan-T5 60M 66.24(0.52) 69.53(0.18) 67.92(0.41) 64.28(0.58) 66.01(0.54) 64.79(0.36)
+OPT 125M 57.02(0.62) 46.53(0.51) 51.25(0.37) 56.87(0.58) 44.32(0.44) 47.89(0.33)
+QWEN-2.5 500M 52.34(0.62) 54.47(0.51) 53.38(0.46) 53.31(0.55) 50.91(0.43) 51.20(0.38)
 PROMPTEHR 75.2M 65.24(0.68) 70.31(0.56) 68.02(0.17) 63.53(0.47) 67.02(0.65) 65.01(0.28)
-LLaMA 7B 68.44(1.03) 69.49(0.70) 69.13(0.31) 67.45(0.93) 66.18(1.12) 66.15(0.56)
+LLaMA-1 7B 68.54(1.12) 69.54(0.73) 69.29(0.32) 67.53(0.91) 66.24(1.13) 66.21(0.64)
 LDAM 41.3M 72.01(0.85) 62.74(0.62) 66.91(0.20) 69.77(0.18) 56.72(0.69) 60.77(0.48)
@@
 EHR-KnowGen 76.9M 65.80(0.64) 70.85(0.45) 68.16(0.11) 63.82(0.53) 67.24(0.55) 65.11(0.13)
-ProMedTS 267.5M 71.63(0.46) 67.81(0.85) 69.69(0.18) 70.12(0.47) 63.58(0.79) 66.21(0.17)
-ProMedTS* 1B 71.12(0.31) 69.33(0.42) 70.21(0.05) 70.97(0.42) 65.51(0.64) 67.56(0.09)
+ProMedTS
+w/ OPT 316M 54.99(0.15) 55.20(0.19) 55.10(0.11) 56.81(0.10) 52.36(0.17) 53.23(0.10)
+w/ QWEN-2.5 685M 48.65(0.25) 72.32(0.17) 58.14(0.16) 57.35(0.20) 66.23(0.13) 57.51(0.12)
+w/ Flan-T5-small 268M 71.63(0.46) 67.81(0.85) 69.69(0.18) 70.12(0.47) 63.58(0.79) 66.21(0.17)
+w/ Flan-T5-large 1B 71.12(0.31) 69.33(0.42) 70.21(0.05) 70.97(0.42) 65.51(0.64) 67.56(0.09)
 Table 1: The performance of comparative methods in the disease diagnosis tasks on MIMIC-III and MIMIC-IV.
@@
 It includes medical discharge summaries, lab test results, chest x-ray images and more.
-Our analysis focuses on EHR data from approximately 27,000 patients including complete medical discharge summaries and lab test results.
+Our analysis focuses on EHR data from approximately 27,000 patients, including complete medical discharge summaries and lab test results.
 The MIMIC-IV dataset Johnson et al.
@@
 Data Pre-processing.
-For medical notes, we extract the brief course from medical discharge summaries, removing numbers, noise, and stopwords.
-Numerical lab test results are converted into time series data using the benchmark tools Harutyunyan et al.
+For medical notes, we extract the brief course from discharge summaries, removing numbers, noise, and stopwords.
+Numerical lab tests are converted into time series data using the benchmark tools Harutyunyan et al.
 (2019b), with missing values filled using the nearest available numbers.
-Time series anomaly descriptions are used the method defined in Appendix A.2.
-Data splitting follows the guidelines in Harutyunyan et al.
-(2019b), using a 4:1 ratio for training and testing.
+Time series anomaly descriptions are used with the method defined in Appendix A.2.
+Data splitting follows the guidelines Harutyunyan et al.
+(2019b) using a 4:1 ratio for training and testing.
 4.2 Baseline Methods
@@
 (2018), DIPOLE Ma et al.
-(2017), PROMPTEHR Wang and Sun (2022), LLaMA–7B Touvron et al.
-(2023), LDAM Niu et al.
-(2021a), FROZEN Tsimpoukelli et al.
+(2017), Flan-T5 Chung et al.
+(2024), OPT Zhang et al.
+(2022), QWEN-2.5 Qwen et al.
+(2025), PROMPTEHR Wang and Sun (2022), LLaMA-1-7B Touvron et al.
+(2023) with anomalies input, LDAM Niu et al.
+(2021), FROZEN Tsimpoukelli et al.
 (2021), and EHR-KnowGen Niu et al.
-(2024).
+(2024a).
 Detailed configurations of these baselines are provided in Appendix A.3.
 For the disease diagnosis task, we adopt two scales of Flan-T5 Chung et al.
-(2024) as the inference LLM to validate our model’s effectiveness in understanding multimodal EHRs, primarily driven by resource considerations and ease of experimentation.
-The Flan-T5-Small-based model is denoted as PromMedTS, while the Flan-T5-Large-based model is denoted as ProMedTS*.
-In principle, any sufficiently LLM could be substituted to potentially achieve even stronger results.
-To ensure a fair comparison, all baselines also employ Flan-T5 as their backbone.
+(2024), OPT-0.1B, and QWEN-2.5-0.5B as the frozen LLM to validate our model’s effectiveness.
+To ensure a fair comparison, all baselines also employ Flan-T5-small as their backbone.
 Reported results are averaged over five runs with different random seeds.
-The statistical significance determined at p < 0.05 by t-test.
-Implementation details for every model are described in Appendix A.4, and training instructions appear in Appendix A.5.
-Our code is publicly available at https://anonymous.4open.science/r/ProMedTS-V1-5F51/.
+The statistical significance was determined at p < 0.05 by t-test.
+Implementation details for every model are described in Appendix A.4, and the example of training instructions appears in Appendix A.5.
+Our code is publicly available at github [^1] https://github.com/Healthcare-Data-Mining-Laboratory/PromptMedTS-V1.
 Models Micro Macro
@@
 4.3 Disease Diagnosis Performance
-Table 1 shows that ProMedTS achieves the highest overall performance, particularly in F1 scores on MIMIC-IV.
-In addition, replacing the LLM with a larger model improves F1 scores on both datasets, indicating our model’s scalability and robustness across different LLMs.
-Furthermore, TimeLLM performs strongly with lab test, highlighting the value of time-series inputs for LLMs in disease diagnosis.
-Text-based methods (e.g., LLaMA) generally outperform time-series approaches, suggesting that medical notes capture richer disease-related information.
-Multimodal models (LDAM, FROZEN, EHR-KnowGen) exceed single-modality baselines (TimeLLM, CAML, DIPOLE, PROMPTEHR), confirming the benefits of integrating text and time series.
-Generative approaches (TimeLLM, LLaMA, FROZEN, EHR-KnowGen) also outperform classification-based methods.
+Table 1 shows the evaluation performance of disease diagnosis for our model ProMedTS with all baselines.
+First, for a single modality model, TimeLLM performs strongly with lab tests, highlighting the value of time-series inputs for LLMs in disease diagnosis.
+Text-based methods (e.g., Flan-T5) generally outperform time-series approaches, suggesting that medical notes capture richer disease-related information.
+Multimodal models (e.g., EHR-KnowGen, LLaMA) exceed single-modality baselines (e.g., TimeLLM, PROMPTEHR), confirming the benefits of integrating text and time series.
+Generative approaches (e.g., TimeLLM, LLaMA, EHR-KnowGen) also outperform classification-based methods.
 Although LLaMA performs well, its higher variance and parameter requirements reduce practicality.
-Notably, the Flan-T5-based ProMedTS and ProMedTS* surpass all baselines in F1 on both datasets, highlighting its efficiency and effectiveness.
+Notably, the performance of our ProMedTS model, which employs Flan-T5-small as its backbone LLM, exceeds that of all baseline methods (especially LLaMA-1-7B) across both micro and macro F1 scores.
+This outcome emphasizes the efficiency and effectiveness of our approach in enabling LLMs to achieve multimodal understanding for disease diagnosis tasks.
+Additionally, ProMedTS consistently improves disease diagnosis performance across various backbone LLMs (Flan-T5-small&large, OPT, QWEN), achieving an average F1 score improvement of around 3%.
+These findings underscore our model’s scalability, robustness, and the generalizability of our approach.
 4.4 Ablation Studies
@@
 ProMedTS w/o ANOMALY removes alignment with anomaly descriptions while keeping alignment between lab test and medical notes to assess the impact of self-supervision.
-Table 2 summarizes the results, showing that ProMedTS w/o LAB suffers a significant drop in F1 scores, highlighting the importance of lab test.
+Table 2 summarizes the results, showing that ProMedTS w/o LAB suffers a significant drop in F1 scores, highlighting the importance of the lab test.
 ProMedTS w/o ANOMALY also shows reduced performance, highlighting the challenges of aligning modalities from discrete and continuous encoding spaces and the adverse effects of misalignment on multimodal understanding.
+These findings emphasize the crucial role of both lab tests and anomaly captions as input for learning accurate time series embeddings, which in turn assist LLMs in disease diagnosis.
 4.4.2 Impact of Self-Supervised Loss Functions
@@
 Both ProMedTS w/o CONTRAST and ProMedTS w/o MATCH show slight declines in F1 scores, emphasizing the importance of $\mathcal{L}_{contrast}$ for aligning and unifying time series and textual inputs within a shared latent space.
-The results also underscore the role of $\mathcal{L}_{match}$ in intra-modal alignment, ensuring the distinctiveness of time series data by aligning lab test with time series prompt embeddings.
+The results also underscore the role of $\mathcal{L}_{match}$ in intra-modal alignment, ensuring the distinctiveness of time series data by aligning lab tests with time series prompt embeddings.
 Notably, ProMedTS w/o GEN exhibits a significant drop in F1 scores, highlighting the critical role of $\mathcal{L}_{gen}$ in refining prompt embeddings and integrating temporal information from time series data and anomaly descriptions.
@@
 Our model, ProMedTS, matches the parameter counts and computation times of multimodal baselines such as LDAM and FROZEN, while using 25× fewer parameters and requiring one-third less training time than LLaMA, all while achieving superior diagnostic performance, highlighting its efficiency and effectiveness in language-time series multimodal alignment and fusion.
+In addition, we conducted an efficiency evaluation using EHR-KnowGen with different backbone LLMs.
+As shown in Table 4, ProMedTS outperformed EHR-KnowGen by an average of 2% improvement in F1 scores while maintaining comparable training times with backbone LLMs of OPT and QWEN-2.5.
+This result supports the effectiveness and efficiency of our model in disease diagnosis.
+4.6 Sensitivity Analysis of Time Series Prompt Length
+We performed a sensitivity analysis to examine the impact of the time series prompt embedding length ($N_{p}$) on the performance of ProMedTS in disease diagnosis.
+Table 5 shows the F1 scores for embedding lengths of 12, 24, and 36.
+Slight fluctuations are observed in both micro and macro F1 scores across datasets.
+The optimal embedding length is 24 for both datasets, consistent with the configuration used in our experiments.
+Models Training Micro F1 Macro F1
+Time (h)
+MIMIC-III
+EHR-KnowGen
+w/ OPT 8h 53.11 52.85
+w/ QWEN-2.5 12h 54.09 51.45
+ProMedTS
+w/ OPT 8h 54.74 53.68
+w/ QWEN-2.5 14h 57.82 55.97
+MIMIC-III
+EHR-KnowGen
+w/ OPT 8.5h 52.46 50.19
+w/ QWEN-2.5 13h 54.19 52.77
+ProMedTS
+w/ OPT 9h 55.10 53.23
+w/ QWEN-2.5 14.5h 58.14 57.51
+Table 4: Performance and Efficiency Comparison on Disease Diagnosis Tasks with Different LLMs.
       $N_{p}$       Micro F1       Macro F1
@@
       36       69.41(0.19)       65.91(0.20)
-Table 4: Sensitivity analysis on different length of time series prompt embedding.
-4.6 Sensitivity Analysis of Time Series Prompt Length
-We performed a sensitivity analysis to examine the impact of the time series prompt embedding length ($N_{p}$) on the performance of ProMedTS in disease diagnosis.
-Table 4 shows the F1 scores for embedding lengths of 12, 24, and 36.
-Slight fluctuations are observed in both micro and macro F1 scores across the MIMIC-III and MIMIC-IV datasets.
-The optimal embedding length is 24 for both datasets, consistent with the configuration used in our experiments.
+Table 5: Sensitivity analysis on different lengths of time series prompt embedding.
+4.7 Evaluating the Role of Anomaly Descriptions
+To highlight the advantages of using lab test anomaly captions over raw numerical time series values in LLMs, we evaluate Flan-T5-small with both input types.
+Table 6 presents the evaluation results on the MIMIC-III and MIMIC-IV datasets for disease diagnosis.
+The results show that Flan-T5 achieves over a 2% improvement in Micro F1 score when using anomaly captions, demonstrating that LLMs interpret anomaly captions more effectively than raw numerical values in time series lab test data.
+Additionally, the inclusion of time series prompts underscores the effectiveness of our model, ProMedTS, in capturing both fine-grained and coarse-grained temporal information from lab test results for disease diagnosis.
 Lab Test Input Micro F1 Macro F1
@@
 Time series prompts 40.14(1.05) 28.43(0.91)
-Table 5: Micro and Macro F1 Scores Across Various Lab Test Input Types on Flan-T5 for Disease Diagnosis
-4.7 Evaluating the Role of Anomaly Descriptions
-To highlight the advantages of using lab test anomaly captions over raw numerical time series values in LLMs, we evaluate Flan-T5-small with both input types.
-Table 5 presents the evaluation results on the MIMIC-III and MIMIC-IV datasets for disease diagnosis.
-The results show that Flan-T5 achieves over a 2% improvement in Micro F1 score when using anomaly captions, demonstrating that LLMs interpret anomaly captions more effectively than raw numerical values in time series lab test data.
-Additionally, the inclusion of time series prompts underscores the effectiveness of our model, ProMedTS, in capturing both fine-grained and coarse-grained temporal information from lab test results for disease diagnosis.
+Table 6: Micro and Macro F1 Scores Across Various Lab Test Input Types on the LLM for Disease Diagnosis
 5 Conclusion and Future Work
-In this paper, we introduce ProMedTS, a lightweight and effective modality fusion framework that leverages self-supervised prompt learning for multimodal EHR integration.
+In this paper, we introduce ProMedTS, a lightweight and effective modality fusion framework that leverages self-supervised time series prompt learning for multimodal EHR integration.
 By bridging the modality gap between medical notes and lab test results, ProMedTS enables LLMs to process structured and unstructured medical data more effectively.
-Its three key modules and self-supervised loss functions advance language–time series integration in healthcare, providing a scalable and adaptable approach for real-world clinical applications.
-Evaluation on two EHR datasets demonstrates that ProMedTS significantly outperforms existing models in disease diagnosis, underscoring its potential to enhance clinical decision-making and improve patient care.
-In future work, we plan to extend our approach to larger and more diverse datasets, explore additional LLM architectures, and investigate further improvements in modality alignment techniques.
+Its three key modules and loss functions advance language–time series integration in healthcare, providing a scalable and adaptable approach for real-world clinical applications.
+Evaluation on two real-world EHR datasets demonstrates that ProMedTS significantly outperforms existing models in disease diagnosis, underscoring its potential to enhance LLMs’ clinical decision-making and improve patient care.
+In future work, we plan to extend our approach to apply to larger and more diverse datasets, such as Time-MMD Liu et al.
+(2024), conduct experiments with different clinical tasks, explore additional LLM architectures, and investigate further improvements in modality alignment techniques.
 Limitations
 While this study focuses on modality alignments and their application in downstream tasks, enhancing the explainability of disease diagnosis remains an area for future work, where we plan to incorporate the Chain-of-Thought rationale Wei et al.
+(2022) via knowledge distillation Hsieh et al.
+(2023); Lin et al.
 (2023).
-Additionally, computational constraints required the use of a relatively compact LLM, limiting the amount of clinical text processed at once, which may impact the model’s ability to leverage full medical histories.
-Expanding to more capable models will help address this challenge.
-Furthermore, our study primarily targets higher-level disease phenotypes in the International Classification of Diseases (ICD) codes Slee (1978), which could be expended to more downstream tasks.
-Future work will explore larger models, such as LLaMA Touvron et al.
-(2023) and Mistral Jiang et al.
-(2023), to improve diagnostic granularity and broaden coverage.
+Furthermore, our study primarily targets higher-level disease phenotypes Slee (1978), which could be expanded to more downstream tasks.
 Ethics Statement
@@
 While the datasets utilized in our research, such as MIMIC-III and MIMIC-IV, are publicly accessible and feature de-identified patient data, accessing these datasets still requires passing the CITI examination and applying for the data through PhysioNet.
+Acknowledgments
+This work is supported by Tencent Rhino-Bird Focused Research Program (Value-aligned Credible Large Language Model) and RMGS project (Artificial Intelligence and Big Data Analytics for Social Good), and is also supported by the National Natural Science Foundation of China (No.62432006, No.62276159).
 Appendix A Appendix
-A.1 Lab Test Anomaly Caption
-Time series anomaly descriptions are generated using the IQR method Vinutha et al.
-(2018) to identify anomalies, capturing their timing and polarity (above or below standard values) and describing them with handcrafted templates.
-To caption the lab test anomaly into textual format, we design several text templates to describe the lab test anomalies.
-All templates are illustrated in Table 6.
-A.2 Algorithm
+A.1 Algorithm
 The training procedure to optimize ProMedTS by minimizing the loss defined in Equation (10) is shown in Algorithm 1.
@@
 11:  end while
-A.3 Baseline Models
-If lab test value is not an abnormal value
+A.2 Lab Test Anomaly Caption
+Time series anomaly descriptions are generated using the IQR method Vinutha et al.
+(2018) to identify anomalies, capturing their timing and polarity (above or below standard values) and describing them with handcrafted templates.
+To caption the lab test anomaly in textual format, we design several text templates to describe the lab test anomalies.
+All templates are illustrated in Table 7.
+If lab test value is not an abnormal value:
 {Lab features} is normal all the time.
-If the lab test value is an abnormal value higher than the standard.
+If the lab test value is an abnormal value higher than the standard:
 {Lab features} is higher than normal {number of times} times.
-If the lab test value is an abnormal value lower than the standard.
+If the lab test value is an abnormal value lower than the standard:
 {Lab features} is lower than normal {number of times} times.
-If the lab test value is an abnormal that include both higher and lower than the standard value.
+If the lab test value is an abnormal that include both higher and lower than the standard value:
 {Lab features} is higher than normal {number of times} times and lower than normal{number of
@@
 Table 7: Lab test anomaly caption template.
+Diagnose disease from the following medical notes and lab test:
+Medical Notes:
+Diagnose disease from the following medical notes: woman decompensated etoh cirrhosis
+initially fatigue ascites hydrothorax uti patient complicated hospital transferred icu hypoxemic
+respiratory failure developed hypotension setting bleeding esophageal varices underwent tips
+banding episodes gi bleeding … prevention outpatient electrolyte abnormalities refeeding syndrome…
+contact narrow urine culture positive plan continue day add fosfomycin …
+Lab test Anomaly Descriptions : $\leftarrow$ Only used during multimodal alignment
+diastolic blood pressure is higher than normal one times, fraction inspired oxygen is higher than
+normal forty-six times, glucose is higher than normal one times, heart rate is normal all the time,
+mean blood pressure is normal all the time, oxygen saturation is normal all the time, respiratory rate
+is higher than normal four times, systolic blood pressure is normal all the time, temperature is normal
+all the time, weight is normal all the time, ph is normal all the time
+Lab Test:
+[[73. 1. 120. 87. 73. 94. 24. 105. 37. 81. 7.53 ],
+[ 79. 1. 120. 87. 94. 94. 30. 131. 37. 81. 7.53 ],
+[73. 1. 217. 87. 73. 94. 24. 105. 37. 81. 7.53 ],
+[73. 1. 120. 87. 73. 94. 22. 105. 37. 81. 7.53 ]]
+Diagnosis:
+Diagnosed Results: Acute and unspecified renal failure, Fluid and electrolyte disorders,
+Septicemia (except in labor), Shock, Chronic obstructive pulmonary disease and bronchiectasis,
+Disorders of lipid metabolism, Cardiac dysrhythmias, Congestive heart failure; nonhypertensive,
+Diabetes mellitus with complications, Other liver diseases.
+Table 8: The Example of Training Instruction Template
+Figure 5: Sensitivity analysis of varying ratios in loss function components, showing Micro and Macro F1 scores for downstream tasks.
+A.3 Baseline Models
   • GRU: The Gated Recurrent Unit (GRU) Cho et al. (2014), a variant of recurrent neural networks (RNNs), employs two gates to capture both long-term and short-term temporal features effectively.
@@
 For a fair comparison with more recent language models, its original embedding layer has been replaced with one from T5.
-    Diagnose disease from the following medical notes and lab test:
-    Medical Notes: ms woman signficant pmh atrial fibrillation lung cancer resection congestive heart
-    hypertension worsening diarrhea dysuria hypotensive admitted unit presumed active issues hypotensive
-    pronounced leukocytosis multiple potential sources ct scan peritoneal fluids free fluid started surgery …
-    Lab Test: $<prefix_{1}>,<prefix_{2}>,...,<prefix_{n}>$
-    Diagnosis:
-    Acute and unspecified renal failure, Fluid and electrolyte disorders, Septicemia (except in labor), Shock,
-    Chronic obstructive pulmonary disease and bronchiectasis, Disorders of lipid metabolism,
-    Cardiac dysrhythmias, Congestive heart failure; nonhypertensive, Diabetes mellitus with complications,
-    Other liver diseases.
-    Table 7: Training Instruction Template
+  • Flan-T5: showcased within the scaling instruction-fine-tuning framework for language models Chung et al.
+(2024).
+It benefits from training on a wide array of datasets geared toward tasks like summarization and question answering.
+  • OPT:OPT is a decoder-only model, designed to match the performance of GPT-3 Brown et al.
+(2020) while requiring only 1/7 of its computational resources.
+Trained primarily on around 180 billion tokens, OPT excels in tasks like poetry generation, dialogue, few-shot translation, and programming Zhang et al.
+(2022).
+  • QWEN-2.5: QWEN-2.5 is a cutting-edge language model, expanding its pre-training dataset to 18 trillion tokens to enhance common sense, expert knowledge, and reasoning.
+It undergoes extensive supervised fine-tuning and multi-stage reinforcement learning, achieving top-tier performance in benchmarks for language understanding, reasoning, mathematics, coding, and human preference alignment Qwen et al.
+(2025).
   • PROMPTEHR: PROMPTEHR Wang and Sun (2022) introduces a novel approach in generative models for electronic health records (EHRs), implementing conditional prompt learning.
@@
   • LDAM: LDAM Niu et al.
-(2021a) leverages multimodal inputs, combining laboratory testing results and medical notes for disease risk prediction.
+(2021) leverages multimodal inputs, combining laboratory testing results and medical notes for disease risk prediction.
 It utilizes label embedding to effectively integrate these two modalities.
@@
   • EHR-KnowGen: EHR-KnowGen Niu et al.
-(2024), touted as the state-of-the-art in EHR multimodal learning models, focuses on disease diagnosis generation.
+(2024a), touted as the state-of-the-art in EHR multimodal learning models, focuses on disease diagnosis generation.
 For this study, external domain knowledge is excluded to ensure a fair comparison.
-Figure 5: Sensitivity analysis of varying ratios in loss function components.
 A.4 Implementation Details
-In experiments, we utilized PyTorch framework version 2.0.1, operating on a CUDA 11.7 environment.
+In experiments, we utilized the PyTorch framework version 2.0.1, operating on a CUDA 11.7 environment.
 We employed the AdamW optimizer with a starting learning rate of $1e^{-5}$ and a weight decay parameter of 0.05.
@@
 To standardize the time series data input, we padded all lab test results to a uniform length of 1000 time steps, allowing us to divide the data into 125 patches, with each patch containing 8 time steps.
-The Flan-T5 is fine-tuned on two MIMIC datasets Johnson et al.
+All LLMs are fine-tuned on two MIMIC datasets Johnson et al.
 (2016, 2023) and then frozen for downstream tasks.